# Hits to Higgs: Reconstruction-Free Higgs Classification from Raw LHC Detector Data Using Higgsformers

### v2

@@ Abstract
We present a comparative study of Higgs event classification at the Large Hadron Collider that bypasses the traditional reconstruction chain. As a benchmark, we focus on distinguishing $t\bar{t}H$ from $t\bar{t}$ events with $H\to b\bar{b}$, a particularly challenging task due to their similar final-state topologies. Our pipeline begins with event generation in Pythia8, fast simulation with ACTS/Fatras, and classification directly from raw detector hits.
We show for the first time that a transformer model originally developed for inner tracker hit-to-track assignment can be retrained to classify Higgs events directly from raw hits. For comparison, we reconstruct the same events with Delphes and train object-based classifiers, including multilayer perceptrons and the Particle Transformer. We evaluate both approaches under varying dataset sizes and pileup levels. Although Higgsformer works exclusively with inner tracker hits (i.e., without calorimeter or muon information), it achieves strong performance with an AUC value of 0.792.
@@ 1 Introduction
The Large Hadron Collider (LHC) at CERN produces proton–proton collisions at an unprecedented scale, up to 40 million events per second, generating hundreds of terabytes of online data and tens of petabytes of stored offline data annually. Traditionally, this massive volume is processed through deeply structured reconstruction pipelines. These include particle trajectory fitting, calorimeter clustering, and jet and object identification, which ultimately reduce the raw detector output into a compact set of high-level observables used in physics analyses. While effective, this approach imposes strong inductive biases and may discard low-level information that could be critical for certain tasks.
@@
This process pair provides a realistic and challenging benchmark: they share similar final-state topologies, but differ subtly in object multiplicities and kinematics, particularly in the presence of additional $b$-jets from the Higgs decay. These subtle differences make it an ideal case study for assessing whether low-level hit information can rival or surpass traditional high-level approaches.
We build a complete end-to-end event simulation pipeline based on Pythia8 Sj_strand_2015 event generation, ACTS/Fatras Spannagel:2019actf detector simulation, and a FlashAttention-accelerated dao2023flashattention2 transformer model to classify events directly from hit patterns. A closely related model was previously developed by us as a Trackformer, used for assigning hits to tracks  caron2025 . In this work, we adapt the same basic architecture for Higgs event classification and refer to it as the Higgsformer. It is a lightweight, set-based Transformer designed to operate directly on raw detector hits.
For comparison, we construct a parallel baseline where the same generated events are processed through Delphes deFavereau:2013fsa to reconstruct standard physics objects (jets, leptons, photons, and missing transverse energy). On this object-level representation, we benchmark against multilayer perceptron (MLP) baselines and the Particle Transformer (ParT) Qu:2022mxj . This setup allows us to highlight the novelty of hit-level end-to-end learning with the Higgsformer, while still enabling a comparison to established object-level approaches. However, this comparison should be interpreted with care, since in this prototype study we contrast a full detector setup, which includes the calorimeter, muon system, and inner tracker, with models that operate only on inner tracker data.
We then evaluate both pipelines under varying detector conditions, including datasets with and without additional proton–proton interactions (pileup), to assess robustness in realistic LHC-like environments.
@@
The remainder of this paper is organized as follows. Section 2 reviews related work. Section 3 describes the event generation procedure and the two parallel datasets: the hit-level dataset obtained with ACTS/Fatras and the reconstructed-object dataset obtained with Delphes. Section 4 introduces the machine learning architectures employed in this study, with particular emphasis on our novel Higgsformer model. Section 5 presents the experimental setup and comparative results, covering performance as a function of dataset size, pileup conditions, and learned feature exploration. Finally, Section 6 summarizes our conclusions andand outlines directions for future work.
@@ 2 Related Work
Våge vage2018 examined the performance of Boosted Decision Trees (BDTs) and Neural Networks (NNs) in the context of $t\bar{t}H(b\bar{b})$ classification at ATLAS. Using standard high-level features such as jet multiplicities, b-tag scores, and invariant masses, the study found comparable results between BDTs and fully connected NNs, both achieving area under the ROC curve (AUC) values in the range of 75–77%. However, it also highlighted that neural networks are more sensitive to class imbalance and training biases, especially in low-statistics regions.
@@
In contrast, our work takes a radically different approach. We eliminate the reconstruction step entirely and operate directly on the raw detector hits using transformer-based models. Inspired by recent work on TrackML caron2025 , we hypothesize that deep architectures can learn to extract task-relevant structure directly from hit patterns given enough training data and realistic detector simulation. To the best of our knowledge, this is the first study exploring end-to-end classification of $t\bar{t}$ vs. $t\bar{t}H$ events from hit-level data alone, without using any handcrafted features or reconstructed quantities.
@@ 3.2 Pileup, postprcessing and reconstruction pipeline
We generated datasets with pileup (PU) levels of 0, 5 and 20 interactions per bunch crossing, each with 40,000 events.  [^1] The full dataset pipeline is available at https://github.com/EugeneShalli/hits-gen. This enables evaluation under varying detector occupancy and complexity.
For the baseline comparison, we generate a parallel dataset using the same Pythia8 configuration described in Section 3.1, ensuring identical signal and background definitions, decay channels, and parton-level cuts. This guarantees a fair comparison between the hit-level and object-level approaches. Detector effects are simulated using Delphes 3.5.0 deFavereau:2013fsa with the default ATLAS detector card. The simulation models the calorimeter, tracking, and muon systems, including realistic resolutions, efficiencies, and acceptance cuts. Jets are reconstructed using the anti-$k_{T}$ algorithm with $R=0.4$, and standard ATLAS-like $b$-tagging efficiencies and mis-tag rates are applied. The output consists of reconstructed high-level physics objects:
  • Electrons and muons with four-momenta and isolation variables.
  • Jets with $p_{T}$, $\eta$, $\phi$, energy, and $b$-tag discriminants.
  • Missing transverse energy and its azimuthal angle.
Objects are ordered by decreasing $p_{T}$ and padded to a fixed multiplicity per category to enable batch processing in the machine learning pipeline. Missing entries are zero-padded, and a corresponding mask array is stored for each event.
@@ 4 Machine Learning Models
We introduce Higgsformer, a lightweight set-based Transformer operating directly on raw hit coordinates $(x,y,z)$. Each hit is treated as a token, embedded into a $d$-dimensional space, and processed with $L$ Transformer encoder layers. Masked mean pooling produces an event-level representation, followed by a linear classifier for binary discrimination.
By default, we use the small configuration: $L=2$ layers, $H=2$ attention heads, $d=16$ hidden size. Larger settings up to $L=4$, $H=4$, $d=64$ were also tested. We employ exact FlashAttention dao2022flashattention for memory-efficient self-attention and mixed precision training (AMP) for speed. Dropout (0.3) and weight decay are applied for regularization.
Training uses binary cross-entropy with logits. We found out that logits are background biased despite the balanced dataset, so we incorporate a positive-class weight of 1.5 to balance classes. We use AdamW optimizer, learning rate $10^{-3}$, batch size 64, and early stopping with patience of 100 epochs. Models are trained up to 500 epochs.
Multilayer perceptron (MLP)
As a baseline, we train an MLP Rumelhart:1986gxv on fixed-length feature vectors assembled from Delphes objects. Up to four jets ($j$), four $b$-jets ($b$), and one particle of each of the remaining types (electron, positron, muon, and anti-muon) are selected in descending $p_{T}$ order. Their $(E,p_{T},\eta,\phi)$ values are flattened into a one-dimensional array. We consider two input configurations: (1) all reconstructed objects together with the MET and its azimuthal angle, which effectively captures the response of the full detector; and (2) only $b$-jets, which provides information primarily associated with the inner tracker, while excluding light jets, leptons, photons, and MET.
Features are normalized using statistics from the training set: $E$, $p_{T}$, and MET via Z-score, and $\eta$, $\phi$ via Min–Max scaling. The MLP has three hidden layers (256-128-64 units, ReLU activations) and a sigmoid output. It is trained using Adam optimizer and binary cross-entropy loss, with early stopping based on validation AUC.
@@
To benchmark against a state-of-the-art object-level model, we employ the Particle Transformer Qu:2022mxj for the $t\bar{t}H$ vs. $t\bar{t}$ classification task. The model operates on a sequence of reconstructed objects, using self-attention layers to capture correlations between particles. The detailed network configuration and hyperparameters follow our previous work Builtjes:2022usj . In this study, we use a simplified version of ParT without interaction matrices or coupling terms.
@@ 5 Experiments and Results
Performance at Dataset Sizes (Pileup 0)
We trained Higgsformer-small on PU 0 datasets of 10k, 20k, and 40k events with identical validation/test splits (1k/1k). We evaluate Higgsformer againts the Baseline classifier which uses ‘number of hits per event’ as a seperating feature only. Test ROC AUC increases monotonically with training size for Higgsformer (Fig. 1), consistent with scaling trends observed in Trackformers caron2025 .
Figure 1: ROC AUC curves on Test set for models trained on datasets with increasing size (10k, 20k, 40k).
@@
We train Higgsformer on the 38k dataset with the varying pileup of 0, 5 and 20. Performance degrades with pileup yet remains well above random; logit histograms (Fig. 2) show that the model still learns to separate classes even at PU 20, while an $n_{\text{hits}}$-only baseline weakens dramatically with increasing pileup (Fig. 3).
(a) PU 0 (AUC =0.779=0.779)
(b) PU 5 (AUC =0.731=0.731)
(c) PU 20 (AUC =0.654=0.654)
Figure 2: Classification output (Logit) histograms for pileup levels 0, 5 and 20. Signal $t\overline{t}h$ is shown in red and background $t\overline{t}$ in blue.
@@
Table 1: Effect of augmentation on Test ROC AUC for Higgsformer-small 38k, pileup=0. +Aug shows the improvement due to data augmentation
Dataset size Higgsformer-small + Aug $\boldsymbol{\Delta}$
10k 0.704 0.721 +0.017
20k 0.757 0.764 +0.007
40k 0.779 0.792 +0.013
@@
We compare our hit-level model to several classifiers trained on reconstructed objects. These include two MLPs with varying inputs trained on fixed-length arrays, and a ParT model that operates on variable-length sequence of physics objects derived from Delphes. As shown in Table 2 and Fig. 4, object-based models achieve high AUC values and show signs of saturation with increasing data, while the hit-level model Higgsformer-small continues to improve with scale. Each model is trained on a dataset of fixed size (10k, 20k, or 40k events), balanced between signal and background. From each dataset, $1\,\mathrm{k}$ events are reserved for validation and $1\,\mathrm{k}$ for testing; the remainder is used for training, denoted as $n_{\text{train}}$. The table reports results at three training sizes, denoted $n_{\text{train}}$ (in thousands): 8, 18, and 38.
Table 2: Test AUC as a function of training set size for object-based modules (MLP, ParT) and the hit-based model (Higgsformer-small). Object-based modules are trained on reconstructed physics objects, while the Higgsformer-small operates directly on raw detector hits.
@@
MLPa (full detector) Reco objects 0.960 0.951 0.959
MLPb (only $b$-jets) 0.859 0.834 0.848
ParTc (full detector) Reco objects 0.972
Higgsformer-small (only inner tracker) Raw detector hits 0.704 0.757 0.779
Higgsformer-small (augmented, only inner tracker) 0.721 0.764 0.792
@@
b Includes only 4 $b$-jets; no jets, no leptons, no photons, or MET.
c Includes the full reconstructed particle set (jets, $b$-jets, leptons, photons, and MET).
Figure 4: AUC as a function of training set size for different model architectures. Higgsformer-small shows continues improvement with scale. The horizontal dashed line indicates the baseline performance at $\mathrm{AUC}=0.616$.
@@
A counts-only classifier based on $n_{\text{hits}}$ provides limited discrimination that diminishes with pileup, while Higgsformer maintains a substantial advantage (Figs. 2 and 3, Table 3).
Table 3: AUC of $n_{\text{hits}}$-only baseline vs. Higgsformer-small.
Pileup AUC ($n_{\text{hits}}$ only) AUC (Higgsformer) Gain
0 0.621 0.765 +0.144
5 0.560 0.731 +0.171
20 0.530 0.654 +0.124
@@
Finally, we evaluate the runtime performance of the models. While classical CPU-based tracking typically requires about 1 s per event, Higgsformer-small achieves inference times below 10 ms on an NVIDIA H100 GPU for pileup = 0; representing several orders of magnitude speed-up.

### v3

# Hits to Higgs: Hit-Level Higgs Classification from Raw LHC Detector Data Using Higgsformer

@@ Abstract
We present Higgsformer, a transformer-based architecture that classifies Higgs events at the Large Hadron Collider directly from raw inner tracker hits, bypassing the traditional reconstruction chain of intermediate physics objects. As a benchmark, we focus on distinguishing $t\bar{t}H$ from $t\bar{t}$ events with $H\to b\bar{b}$, a particularly challenging task due to their similar final state topologies. Our pipeline begins with event generation in Pythia8, fast simulation with ACTS/Fatras, and classification directly from raw detector hits.
We show for the first time that a transformer model originally developed for inner tracker hit-to-track assignment can be retrained to classify Higgs signal events directly from raw hits. For comparison, we reconstruct the same events with Delphes and train a Particle Transformer as a object-based classifier. We evaluate both approaches under varying dataset sizes and pileup levels. Despite relying exclusively on inner tracker hits, our large Higgsformer achieves an AUC of $0.855$, matching the performance of the traditional reconstruction pipeline at a b-tagging efficiency of $\approx 40\%$ under the same detector constraints.
@@ 1 Introduction
The Large Hadron Collider (LHC) at CERN produces proton–proton collisions at an unprecedented scale, up to 40 million events per second, generating hundreds of terabytes of online data and tens of petabytes of stored offline data annually. Traditionally, this massive volume is processed through deeply structured reconstruction pipelines. These include particle trajectory fitting, calorimeter clustering, and jet and object identification, which ultimately reduce the raw detector output into a compact set of high-level observables used in physics analyses as shown in Fig. 1. While effective, this approach imposes strong inductive biases and may discard low-level information that could be critical for certain tasks.
@@
Figure 1: Reconstruction pipeline overview, compared to learning from raw detector data. Classical reconstruction pipeline compresses $10^{4}$ raw detector hits per event to $10^{1}$ reconstructed objects.
@@
This process pair provides a realistic and challenging benchmark: they share similar final state topologies, but differ subtly in object multiplicities and kinematics, particularly in the presence of additional $b$-jets from the Higgs decay. These subtle differences make it an ideal case study for assessing whether low-level hit information can rival or surpass traditional high-level approaches.
We build a complete end-to-end event simulation pipeline based on Pythia8 Sj_strand_2015 event generation, ACTS/Fatras Spannagel:2019actf detector simulation, and a FlashAttention-accelerated dao2023flashattention2 transformer model to classify events directly from hit patterns. A closely related model was previously developed by us as a Trackformer, used for assigning hits to tracks  caron2025 . In this work, we adapt the same basic architecture for Higgs event classification and refer to it as the Higgsformer. It is a lightweight, set-based Transformer designed to operate directly on raw detector hits. In addition, we demonstrate the capabilities of a larger model.
For comparison, we construct a parallel baseline in which the same generated events are processed through Delphes deFavereau:2013fsa to reconstruct track-jets, their $b$-tag information, and track-based missing transverse momentum. On this object-level representation, we benchmark against the Particle Transformer (ParT) Qu:2022mxj . This setup allows us to highlight the novelty of hit-level end-to-end learning with the Higgsformer, while still enabling a comparison to established object-level approach. We then evaluate both pipelines under varying detector conditions, including datasets with and without additional proton–proton interactions (pileup), to assess robustness in realistic LHC-like environments.
@@
We emphasize that this work is a simulation-based proof-of-concept and does not establish a complete deployment strategy on collision data. Hit-level classifiers can be more sensitive to residual data–simulation differences than approaches based on standard reconstructed objects with mature calibration/validation programs. A realistic application would therefore require integration into the experiment’s standard workflow, together with robustness studies, calibrations and data-driven tests in control samples and regions. This is beyond the scope of the present paper and left for future work.
The remainder of this paper is organized as follows. Section 2 reviews related work. Section 3 describes the event generation procedure and the two parallel datasets: the hit-level dataset obtained with ACTS/Fatras and the reconstructed-object dataset obtained with Delphes. Section 4 introduces the machine learning architectures employed in this study, with particular emphasis on our novel Higgsformer model. Section 5 presents the experimental setup and comparative results, covering performance as a function of dataset size, pileup conditions, and learned feature exploration. Finally, Section 6 summarizes our conclusions and outlines directions for future work.
@@ 2 Related Work
Våge vage2018 examined the performance of Boosted Decision Trees (BDTs) and Neural Networks (NNs) in the context of $t\bar{t}H(b\bar{b})$ classification at ATLAS. Using standard high-level features such as jet multiplicities, b-tag scores, and invariant masses, the study found comparable results between BDTs and fully connected NNs, with both achieving an area under the receiver operating characteristic (ROC) curve, or AUC, in the range of 75–77%. However, it also highlighted that neural networks are more sensitive to class imbalance and training biases, especially in low-statistics regions.
@@
A complementary line of work uses ML to perform particle- or object-level reconstruction starting from low-level detector information, providing a middle ground between conventional object-based pipelines and direct hit-level classification. Examples include learning-based tracking that reconstructs charged-particle trajectories from hits via learned hit association and track building Ju:2021exatrkx , ML-assisted calorimeter clustering/shower reconstruction in high-granularity calorimeters Bhattacharya:2022hgcal , and ML-based particle-flow reconstruction that predicts a set of particle candidates from tracks and calorimeter clusters Pata:2021mlpf . These approaches explicitly produce reconstructed particles/objects and can therefore interface naturally with established calibration and validation workflows.
In contrast, our work takes a radically different approach. We eliminate the reconstruction step entirely and operate directly on the raw detector hits using transformer-based models. Inspired by recent work on TrackML caron2025 , we hypothesize that deep architectures can learn to extract task-relevant structure directly from hit patterns given enough training data and realistic detector simulation. To the best of our knowledge, this is the first study exploring end-to-end event level classification at LHC experiment, in particular $t\bar{t}$ vs. $t\bar{t}H$ events, from hit-level data alone, without using any handcrafted features or reconstructed quantities.
@@ 3.2 Pileup, postprcessing and reconstruction pipeline
We generate two different datasets: hit-level dataset and reconstructed objects-level dataset. Generation pipeline is shown in Fig. 2.
Figure 2: Pipeline for hit-level data generation with Open Data Detector and ACTS (up), and reconstructed objects-level data generation with Delphes (bottom).
@@
We generated datasets with pileup (PU) levels of 0, 5 and 20 interactions per bunch crossing, each with 40,000 events.  [^1] The full dataset pipeline is available at Hits-Gen. The dataset for pileup 0 is available at Dataset Pileup 0. This enables evaluation under varying detector occupancy and complexity.
For the object-level comparison we simulate detector effects with Delphes 3.5.0 deFavereau:2013fsa using the default ATLAS card. We define an inner tracker-only baseline to better match the information content available to Higgsformer. In this tracker-only setup we restrict the inputs to track-based objects: track-jets reconstructed with the TrackJetFinder module and track-based missing transverse momentum from TrackMissingET; reconstructed leptons and photons are not included since they rely on calorimeter and/or muon-system information. Track-jet $b$-tagging follows the ATLAS-card implementation via the TrackBTagging module, which assigns a tag bit to each track-jet using a $p_{T}$-dependent efficiency parameterization for $b$-jets (flavour index {5} in the card),
$\epsilon_{b}(p_{T})=f\,\tanh(0.003\,p_{T})\left(\frac{30}{1+0.086\,p_{T}}\right),$ (1)
where $f$ is an overall prefactor that approximately sets the working point. We assess the sensitivity of the object-level baselines to the assumed tagging performance by varying only this $b$-jet working point, considering f=80%, 60%, 40%. [^2] This Delphes implementation is a simplified, fast-simulation parameterization meant to emulate typical fixed-efficiency $b$-tagging operating points. In ATLAS analyses, $b$-tagging is performed with multivariate taggers and the corresponding efficiencies are measured in collision data and applied to simulation via data-to-simulation scale factors as functions of jet kinematics and event conditions; see e.g. Ref. ATLAS:2019btag . Track-jet $b$-tagging in boosted topologies (track-jets matched to large-$R$ calorimeter jets) is used and validated in ATLAS as well; see e.g. Ref. ATLAS:2019boostedHbb .
The output consists of reconstructed high-level physics objects
  • Track-jets with four-momenta $(p_{T},\eta,\phi,E)$ and $b$-tag discriminants.
[…]
Objects are ordered by decreasing $p_{T}$ and padded to a fixed multiplicity per category to enable batch processing in the ML pipeline. Missing entries are zero-padded, and a corresponding mask array is stored for each event.
@@ 4 Machine Learning Models
We evaluate two different models: Higgsformer and ParT, trained and operating on a raw detector-level data and reconstructed objects data respectively as shown in Fig. 3.
Figure 3: ParT (up) trained on reconstructed objects ($10^{1}$ inputs per event), and Higgsformer (bottom) trained on detector hits ($10^{4}$ inputs per event).
Higgsformer
We introduce Higgsformer, a lightweight set-based Transformer operating directly on raw hit coordinates $(x,y,z)$. Each hit is treated as a token, embedded into a $d$-dimensional space, and processed with $L$ Transformer encoder layers. We present two models: Higgsformer-small and Higgsformer-big.
Higgsformer-small uses $L=2$ encoder layers, $H=4$ attention heads, and hidden size $d=32$. Input features are projected with a linear layer followed by LayerNorm. The encoded hit representations are aggregated with masked mean pooling to obtain an event-level representation, which is passed to a linear classifier for binary signal-background discrimination. For efficiency, Higgsformer-small uses FlashAttention dao2022flashattention and mixed precision training (AMP).
Higgsformer-big uses the EncCla Trackformer encoder backbone caron2025 adapted to multi-task learning. It employs $L=8$ layers, $H=8$ attention heads, and hidden size $d=128$. As in the Higgsformer-small model, hit features are first linearly embedded and then processed with Transformer encoder layers. For the classification branch, several aggregation strategies were studied, namely mean pooling, max pooling, and a learnable CLS token. No significant performance differences were observed between the global pooling strategies, and max pooling was used as the default choice. In addition to the binary classification head, Higgsformer-big includes an auxiliary $H_{T}$ regression head implemented as a two-layer MLP on the max-pooled event representation. This auxiliary objective is used to encourage the model to learn more physically meaningful event-level features. For memory-efficient attention on long sequences, Higgsformer-big uses FlexAttention dong2024flex with dynamic padding masks, and can be initialised from pretrained Trackfromer encoder weights in the future.
Training uses binary cross-entropy for classification and AdamW optimizer. Higgsformer-small is trained with learning rate $10^{-4}$ for 30 epochs with batch size 128, ReduceLROnPlateau and early stopping with patience of 10 epochs. Higgsformer-big is trained with binary cross-entropy and Huber loss for $H_{T}$ regression with equal weight contributions in the loss. Higgsformer-big is trained with learning rate $5*10^{-5}$, the other settings are unchanged. Dropout (0.2) and weight decay are applied for regularization for both models.
@@
To benchmark against a state-of-the-art object-level model, we employ the Particle Transformer Qu:2022mxj for the $t\bar{t}H$ vs. $t\bar{t}$ classification task. The model operates on a sequence of reconstructed objects, using self-attention layers to capture correlations between particles. The detailed network configuration and hyperparameters follow our previous work Builtjes:2022usj . In Ref. Builtjes:2022usj , several physics-informed ParT variants were developed in which pairwise kinematic features and Standard Model interaction matrices were incorporated into the attention mechanism. In the present work, these extensions are intentionally omitted and the model is used in its vanilla self-attention form, in order to provide a clean object-level baseline for comparison with the hit-level Higgsformer approach.
@@ 5 Experiments and Results
Performance at Dataset Sizes (Pileup 0) and Model Scale
We trained Higgsformer-small on PU 0 datasets of 10k, 20k, and 40k events with identical validation/test splits (1k/1k). We evaluate Higgsformer against the Baseline classifier which uses ‘number of hits per event’ as a separating feature only. Test ROC AUC increases monotonically with training size for Higgsformer (Fig. 4), consistent with scaling trends observed in Trackformers caron2025 . Higgsformer-big yields a substantial further improvement over Higgsformer-small, demonstrating the benefit of model scale.
Figure 4: ROC AUC curves on Test set for Higgsformer-small models trained on datasets with increasing size (10k, 20k, 40k), and Higgsformer-big model trained on 40k dataset with augmentation.
@@
We train 3 different Higgsformer-small models on the 38k dataset generated with the varying pileup of 0, 5 and 20, and evaluate those models on the 2k test set of respective pileup level. Performance degrades with pileup yet remains well above random; logit histograms (Fig. 5) show that the model still learns to separate classes even at PU 20, while an $n_{\text{hits}}$-only baseline weakens dramatically with increasing pileup (Fig. 6). Additionally we show logit histograms for Higgsformer-big trained on 38k pileup 0 dataset to show that it learns better class separation.
(a) PU 0 Higgsformer-small (AUC =0.773=0.773)
(b) PU 0 Higgsformer-big (AUC =0.856=0.856)
(c) PU 5 Higgsformer-small (AUC =0.731=0.731)
(d) PU 20 Higgsformer-small (AUC =0.654=0.654)
Figure 5: Classification output (Logit) histograms for pileup 0 for Higgsformer-small (a) and Higgsformer-big (b) models, and pileup 5 (c) and 20 (d) for Higgsformer-small. Signal $t\overline{t}H$ is shown in red and background $t\overline{t}$ in blue.
@@
Table 1: Effect of geometric data augmentation on the test ROC AUC of Higgsformer-small at zero pileup for different training set sizes. The column ‘+ Aug’ reports the AUC obtained with augmentation, while $\Delta$ denotes the improvement over the non-augmented model.
Dataset size Higgsformer-small + Aug $\boldsymbol{\Delta}$
10k 0.704 0.723 +0.019
20k 0.745 0.761 +0.016
40k 0.773 0.791 +0.018
@@
We compare our hit-level models against several classifiers trained on reconstructed objects using only inner tracker information. The ParT models operate on variable-length sequences of physics objects derived from Delphes and are evaluated at three b-tagging working points (WPs) corresponding to efficiencies of 0.4, 0.6, and 0.8. As shown in Table 2 and Fig. 7, all object-based models achieve high AUC values and exhibit signs of saturation with increasing training data. In contrast, Higgsformer-small and Higgsformer-big continue to improve with scale. Remarkably, Higgsformer-big reaches an AUC comparable to the object-based model at the $f=0.4$ working point.
Figure 7: Test AUC as a function of training set size for ParT trained on reconstructed objects at three b-tagging WPs and for the augmented Higgsformer models trained on raw inner tracker hits. While the ParT performance is close to saturation, the augmented Higgsformer models continue to improve with scale. At 38k training events, Higgsformer-big reaches an AUC comparable to ParT at the $f=0.4$ b-tagging WP. The horizontal dashed line shows the counts-only baseline ($\mathrm{AUC}=0.616$).
Table 2: Test AUC as a function of training set size for object-based module (ParT) and the hit-based models (Higgsformer-small and Higgsformer-big). Object-based module is trained on reconstructed physics objects, while the Higgsformer operates directly on raw detector hits.
@@
ParTa (WP = 80%) Reco objects 0.949(4) 0.951(4) 0.952(2)
ParTa (WP = 60%) Reco objects 0.865(2) 0.867(5) 0.872(5)
ParTa (WP = 40%) Reco objects 0.842(8) 0.849(8) 0.852(14)
Higgsformer-bigb (augmented, only inner tracker) Raw detector hits 0.820(9) 0.840(32) 0.855(16)
Higgsformer-small (augmented, only inner tracker) 0.723(98) 0.761(20) 0.791(12)
@@
b Includes $H_{T}$ regression
@@
A counts-only classifier based on $n_{\text{hits}}$ provides limited discrimination that diminishes with pileup, while Higgsformer maintains a substantial advantage (Figs. 5 and 6).
@@
a avg(H) – average importance of Higgs-related hits in an event.
b avg(non-H) – average importance of hits in an event, originated from other particles.
[…]
@@
Finally, we evaluate the runtime performance of the models. While classical CPU-based tracking typically requires about 1 s per event, Higgsformer-small and Higgsformer-big inference times are below 2 ms and 10 ms respectively on an NVIDIA A100 GPU for pileup = 0; representing several orders of magnitude speed-up.